OpenStudy (anonymous):

if the remainder of the quotient a/b is 0, then the decimal form of a/b is a ____

myininaya (myininaya):

\[\frac{a}{b}=Q +\frac{R}{b}=Q+\frac{0}{b}=Q+0=Q \in \mathbb{Z}\]

OpenStudy (anonymous):

What?????

myininaya (myininaya):

for example \[\frac{20}{5} \text{ has 0 as a remainder} \] \[\frac{20}{5} =4+\frac{0}{5}=4+0=4 \in \mathbb{Z}\]

OpenStudy (anonymous):

what is the e and z

myininaya (myininaya):

\[\mathbb{Z} =\text{ the set of integers} , \in \text{ means element of }\]

myininaya (myininaya):

4 is an element of the set of integers

OpenStudy (anonymous):

im so con

OpenStudy (anonymous):

*confused

myininaya (myininaya):

confused about a/b being an element of the integers?